Annual maximum wind speed gusts have been modeled as both a Type I (Gumbel) and Type II largest extreme value. If the average maximum wind gust is 60 mph and the standard deviation is 12 mph, what is the probability that the annual maximum wind speed gust will exceed 90 mph for each model?
Answer #1

mean = 60


std = 12


X = 90


P(X>90) = P(Z > X -mean/std)


= P(Z > 90 - 60/12)


= P(Z > 2.5)


=0.0062


0r


0.62%